OpenStudy (juana02):

Topsoil A 40-pound bag of topsoil will cover a surface area of 12 ft2. a) How many pounds of topsoil are needed to cover a surface area of 480 ft2 ? b) How many bag of topsoil must be purchased to cover a surface area of 480 ft2 ?

OpenStudy (danjs):

\[\frac{ 480~ft^2 }{ 1 }*\frac{ 40~lb }{ 12~ft^2 } = ? ~lb\]

OpenStudy (danjs):

ft^2 cancels, you are left with pounds of soil

OpenStudy (danjs):

Then use that answer from part a) in this for part b \[\frac{ ?~pounds }{ 1 } * \frac{ 1~Bag }{ 40~lb } =~~ ? ~~Bag\]

OpenStudy (juana02):

i got for a 1600

OpenStudy (juana02):

1600 pounds

OpenStudy (danjs):

right

OpenStudy (juana02):

and for b is 40

OpenStudy (danjs):

You see how i set it up, you just start with what you are given, then multiply by ratios of 1 to cancel units in your favor, Like in part a, you start with 480 ft^2 Then multiply it by a ratio of 1, 40 lb = 12 ft^2 , you want the f^2 to cancel, so you put it in the bottom

OpenStudy (danjs):

Then for part b , you use 40 lb = 1 bag

OpenStudy (juana02):

yes i want to know how u set it up from the word problem

OpenStudy (danjs):

You can always multiply by fractions that equal 1, like \[\frac{ 1~bag }{ 40~lb }\] they told you 1 bag is 40 lb

OpenStudy (danjs):

worry about the units first, arrange the problem so the units cancel out , then take care of the numbers

OpenStudy (danjs):

Given: 1600 lb, convert to bags \[\frac{ 1600~lb }{ 1 } * \frac{ 1~Bag }{ 40~lb } =~~ ? ~~Bag\]

OpenStudy (danjs):

1 bag = 40 lb, put lb in the bottom , so lb cancels from top and bottom, you are left with bags

OpenStudy (danjs):

Just remember to get the units to work out first

OpenStudy (danjs):

that make it a bit easier?

OpenStudy (juana02):

yes